Find the exact value of the six trigonometric functions of θ if the terminal side of θ in standard position contains the point (-5,-4).
Please Help.
Thanks

Answer by stanbon(48564)

Find the exact value of the six trigonometric functions of θ if the terminal side of θ in standard position contains the point (-5,-4).
x=-5,y=-4, find r as follows:
r=sqrt(5^2 + 4^2) = sqrt(41)
Then:
sine = y/r
cosine = x/r
tangent = y/x
csc = r/y
sec = r/x
cot = x/y
You can make the number substitutions.
Cheers,
Stan H.